1. An airplane flying at 50.0 m/s is bringing food and emergency first aid supplies to a camp. The plan is for the pilot to drop the supplies so that they land on an “X” marked on the ground 150 m below.

a)How long will it take the supplies to fall to the ground? (You can ignore the effect of air resistance.)

b)How far in front of the “X” should the pilot release the supplies so that they land directly on the “X”?



1
Expert's answer
2021-02-25T11:24:04-0500

a)


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(150)}{9.8}}=5.5\ s"

b)


"x=vt=(50)(5.5)=275\ m"